OpenStudy (anonymous):

Solve and show work 15n+16+4n=130 x=?

OpenStudy (anonymous):

i mean what does h=?

OpenStudy (anonymous):

you want this one or that new one?

OpenStudy (anonymous):

both :)

OpenStudy (anonymous):

so. . . . . . . . . .

OpenStudy (anonymous):

oh sorry, I wasn't paying attention

OpenStudy (anonymous):

oh thx -____-

OpenStudy (anonymous):

15n+16+4n=130 15n+4n+16-16=130-16 19n=114 n=114/19 clear?

OpenStudy (anonymous):

and someone have already done the other for you

OpenStudy (anonymous):

n= wat

OpenStudy (anonymous):

pay attention!!

OpenStudy (anonymous):

n=114/19 oh yeah, lol, n=6. 114/19=6 sorry haha

OpenStudy (anonymous):

lol thx :D

OpenStudy (anonymous):

np